OpenStudy (anonymous):

What is the discriminant of x2 – 7x + 2 = 0? 41 –57 0 –1

OpenStudy (cwrw238):

(-7)^2 - 4*1*2

OpenStudy (cwrw238):

= 49 - 8

OpenStudy (ajprincess):

For a quadratic eqn ax^2+bx+c=0 discriminant=b^2-4ac